OpenStudy (anonymous):

A function has the function rule y = x - 7. If the input is 4, what is the output?

OpenStudy (anonymous):

y = f(x) The input is 4, so "put" 4 into f(x). See what I did there? f(4) = (4) - 7 = -3 The output is -3. Is this clear?

OpenStudy (anonymous):

yes and thank you

OpenStudy (anonymous):

Glad I helped.